### v2

# Exploring the unleaved tree of numerical semigroups up to a given genus

Abstract
[…]
1 Introduction
A numerical semigroup is a subset $\Lambda$ of ${\mathbb{N}}$ that contains 00, is closed under addition, and has a finite complement in ${\mathbb{N}}_{0}$. The elements in ${\mathbb{N}}_{0}\setminus\Lambda$ are called the gaps of the semigroup, the largest gap is called the Frobenius number, and the number of gaps is the genus $g(\Lambda)$ of the numerical semigroup.
There have been many efforts to compute the sequence $n_{g}$ counting the number of numerical semigroups of genus $g$ and today the sequence values are known up to $n_{75}$ [14, 4, 12, 9, 11]. See [19] (https://oeis.org/A007323) for the complete list and for more information. It was conjectured in 2007 that the sequence $n_{g}$ is increasing, that each term is at least the sum of the two previous terms, and that the ratio between each term and the sum of the two previous terms approaches one as $g$ grows to infinity, which is equivalent to have a growth rate approaching the golden ratio $\frac{1+\sqrt{5}}{2}$ [3, 4]. The last statement of the conjecture was proved by Alex Zhai [20].
A numerical semigroup $\Lambda$ is generated by a set of integers $a_{1},\dots,a_{k}$ if $\Lambda=a_{1}{\mathbb{N}}_{0}+\dots+a_{k}{\mathbb{N}}_{0}$. In this case $a_{1},\dots,a_{k}$ need to be coprime and we write $\Lambda=\langle a_{1},\dots,a_{k}\rangle$. Each numerical semigroup has a unique minimal set of generators. The set of elements smaller than the Frobenius number are called the left elements. The minimal generators that are not left elements are called right generators. The elements of $\Lambda$ can be enumerated in increasing order as $\lambda_{0}=0,\lambda_{1},\dots$.
Special cases of numerical semigroups are the ordinary semigroups, whose gaps are all consecutive from the integer $1$, quasi-ordinary semigroups, whose gaps are all consecutive from the integer $1$, except for an isolated gap, indeed, the Frobenius number, and pseudo-ordinary semigroups, which are the union of an ordinary semigroup and an element not in the ordinary semigroup.
The problem of defining an algorithm to either count or visit all numerical semigroups up to a given genus has commonly been tackled by means of the so-called tree of numerical semigroups ${\mathscr{T}}$. This tree has the trivial semigroup ${\mathbb{N}}_{0}$ as its root and the children of each node are the semigroups obtained by taking away one by one the right generators of the parent. This construction was first considered in [16, 18, 17].
Given a numerical semigroup $\Lambda$, consider the conductor $c(\Lambda)$, which is the first non-gap larger than the Frobenius number, the multiplicity $m(\Lambda)$, which is its first non-zero non-gap, the jump $u(\Lambda)$, which is the difference between the second and the first non-zero non-gaps, and the efficacy $r(\Lambda)$, which is the number of right generators. An encoding of a numerical semigroup will be a finite set of finite parameters uniquely defining the numerical semigroup, together with its conductor, miltiplicity, jump, efficacy and genus. We will describe a general framework for algorithms exploring the semigroup tree up to a given genus for a general encoding. This framework fits the previous RGD-algorithm [9] and the seeds algorithm [8, 6]. It just involves the notions of right generators, strong generators (defined in Section 2), pseudo-ordinary and quasi-ordinary semigroups.
Then we will present an encoding system, that, used in the previous farmework, will allow to trim all the leaves and branches of ${\mathscr{T}}$ not arriving to a given depth, providing the so-called unleaved tree of numerical semigroups of a given depth (see Figure 1). This represents a new paradigm in the exploration of the semigroup tree up to a given genus, which will lead to much more efficient algorithms than the ones used up to date. It has been already used to compute the number of semigroups of genus $76$ and $77$:
[…]
2 General algorithms for visiting semigroups up to a given genus
Suppose that we can associate a unique finite set of finite parameters $E(\Lambda)$ to each numerical semigroup $\Lambda$ such that, the set $E(\Lambda)$ together with $g(\Lambda)$, $c(\Lambda)$, $m(\Lambda)$, $u(\Lambda)$, $r(\Lambda)$, completely determines $\Lambda$. We call it an encoding of $\Lambda$. Examples of encodings are
  • The minimal set of generators;
  • The Apéry set, defined as the minimal elements of $\Lambda$ of each congruence class modulo $m$, or the set of Kunz coordinates, defined as the integer quotients of the elements in the Aréry set when divided by $m$;
  • The set of left elements or the set of gaps;
  • The sequence $\nu_{i}=\#\{\lambda_{j}\in\Lambda:\lambda_{i}-\lambda_{j}\in\Lambda\}$ up to $i=2c-g-1$ [13, 1];
  • The $\oplus$ operation defined by $i\oplus j=k$ if $\lambda_{i}+\lambda_{j}=\lambda_{k}$ for all integers $i,j$ with $10\leqslant i,j\leqslant 2c-g-$ 1 (indeed, $\nu_{i}=\#\{(j,k)\in{\mathbb{N}}^{2}:j\oplus k=i\}$) [2];
  • The decomposition numbers $d_{i}=\lceil\frac{\nu_{i}}{2}\rceil$ up to $i=2c-g-1$ [12];
  • The right-generators descendant (RGD) of $\Lambda$, defined as the numerical semigroup obtained by removing from $\Lambda$ all its right generators, up to its $(c+m)$th element [9];
  • The bitstreams encoding the gaps and the set of seeds of order up to $c-g$, where a $p$th order seed is an element larger than the Frobenius number and necessarily smaller than $c+\lambda_{p}-\lambda_{p-1}$, such that $\lambda_{s}+\lambda_{p}\neq\lambda_{i}+\lambda_{j}$ for all $p<i\leqslant j<s$ [8, 6].
One can prove that all right generators of a semigroup $\Lambda$ with conductor $c$ and multiplicity $m$ are at most $c+m-1$. Furthermore, it was already proved in [5] that if $\sigma_{1},\dots,\sigma_{r}$ are the right generators of $\Lambda$, with $\sigma_{1}<\sigma_{2}<\dots<\sigma_{r}$, then the right generators of $\Lambda\setminus\{\sigma_{i}\}$ are either $\{\sigma_{i+1},\dots,\sigma_{r}\}$ or $\{\sigma_{i+1},\dots,\sigma_{r},\sigma_{i}+m\}$. We say that $\sigma_{i}$ is a strong generator of $\Lambda$ in the second case. Otherwise, we say that it is a weak generator. We will call $\sigma_{1}$ the primogenial right generator of $\Lambda$ and $\Lambda\setminus\{\sigma_{1}\}$ the primogenial child of $\Lambda$. If $i\neq 1$, we say that $\sigma_{i-1}$ is the predecessor sibling of $\sigma_{i}$ and $\Lambda\setminus\{\sigma_{i-1}\}$ is the predecessor sibling of $\Lambda\setminus\{\sigma_{i}\}$.
Hence, for the exploration of ${\mathscr{T}}$ one is interested in encodings of semigroups for which: there exists an efficient procedure to identify right generators and strong right generators; there exists an efficient procedure to obtain $E(\Lambda\setminus\{\sigma_{i}\})$, either from the encoding of the parent $E(\Lambda)$ or from the encoding of the predecessor sibling $E(\Lambda\setminus\{\sigma_{i-1}\})$, should it exist (i.e., if $i\neq 1$).
Furthermore, we have the following results from [9].
Lemma 2.1.
[9, Lemma 3.1 and Lemma 3.3] Let $\Lambda$ be a non-ordinary semigroup with multiplicity $m$ and jump $u$. If $\sigma\geqslant c+u$, then $\sigma$ is not a strong generator.
Let $\Lambda$ be a pseudo-ordinary semigroup with multiplicity $m$ and jump $u$. Then $c=m+u$ and all integers between $c$ and $c+m-1$ are right generators except the integer $2m$. Furthermore, a right generator $\sigma$ is strong if and only if $\sigma<c+u$.
Suppose we have the following procedures
  • CheckRightGenerator($E(\Lambda),g,c,m,u,r,\sigma$)$=$ CRG($E(\Lambda),g,c,m,u,r,\sigma$), to check whether $\sigma$ is a right generator of $\Lambda$;
  • CheckStrongGenerator($E(\Lambda),g,c,m,u,r,\sigma$)$=$ CSG($E(\Lambda),g,c,m,u,r,\sigma$), to check whether a right generator $\sigma$ is a strong generator of $\Lambda$;
  • EncodingFromParent($E(\Lambda),g,c,m,u,r,\sigma$)$=$ EFP($E(\Lambda),g,c,m,u,r,\sigma$), to obtain $E(\Lambda\setminus\{\sigma\})$, where $\sigma$ is a right generator of $\Lambda$;
  • For $i\neq 1$, EncodingFromPredecessorSibling($E(\Lambda\setminus\{\sigma_{i-1}\}),g,c,m,u,r,\sigma_{i}$)$=$ EFPS($E(\Lambda\setminus\{\sigma_{i-1}\}),g,c,m,u,r,\sigma_{i}$), to obtain $E(\Lambda\setminus\{\sigma_{i}\})$, where $\sigma_{i}$ is a non-primogenial right generator and $\sigma_{i-1}$ is its predecessor sibling.
Using these procedures and Lemma 2.1, we will define a general algorithm for visiting semigroups of a given genus. We first present two subprocedures and at the end the general procedure.
The procedure Descend detailed in Algorithm 1 explores recursively the descendants of a non-ordinary and non-pseudo-ordinary semigroup up to genus $\gamma$. Its input parameters are $E(\Lambda)$, $m=m(\Lambda)$, $u=u(\Lambda)$, $c=c(\Lambda)$, $g=g(\Lambda)$, $r=r(\Lambda)$, and $\gamma$.
Algorithm 1 Descend
procedure Descend($E(\Lambda),g,c,m,u,r,\gamma$)
     Visit $\Lambda$
     if $g<\gamma$ then
         $\tilde{r}\leftarrow r$
         $\tilde{E}\leftarrow E$
         for $\sigma$ from $c$ to $c+u-1$ $\triangleright$ where strong generators may occur
              if CRG($\tilde{E},g,c,m,u,\tilde{r},\sigma$) $=$ true  then
                  if $\sigma$ is primogenial then
                       $\tilde{E}\leftarrow$ EFP($\tilde{E},g,c,m,u,\tilde{r},\sigma$)
                  else
                       $\tilde{E}\leftarrow$ EFPS($\tilde{E},g,c,m,u,\tilde{r},\sigma$)
                  end if
                  if CSG($\tilde{E},g,c,m,u,r,\sigma$) $=$ true  then
                       Descend($\tilde{E},g+1,\sigma+1,m,u,\tilde{r},\gamma$)
                       $\tilde{r}\leftarrow\tilde{r}-1$
                  else
                       $\tilde{r}\leftarrow\tilde{r}-1$
                       Descend($\tilde{E},g+1,\sigma+1,m,u,\tilde{r},\gamma$)
                  end if
              end if
         end for
         while $\tilde{r}>0$ do $\triangleright$ no more strong generators
              if CRG($\tilde{E},g,c,m,u,\tilde{r},\sigma$) $=$ true  then
                  if $\sigma$ is primogenial then
                       $\tilde{E}\leftarrow$ EFP($\tilde{E},g,c,m,u,r,\sigma$)
                  else
                       $\tilde{E}\leftarrow$ EFPS($\tilde{E},g,c,m,u,r,\sigma$)
                  end if
                  $\tilde{r}\leftarrow\tilde{r}-1$
                  Descend($\tilde{E},g+1,\sigma+1,m,u,\tilde{r},\gamma$)
              end if
         end while
     end if
end procedure
The procedure PseudoDescend explores the children of a given pseudo-ordinary semigroup $\Lambda$ that is not pseudo-ordinary, and calls Descend to explore all their respective descendants in ${\mathscr{T}}$. It is detailed in Algorithm 2.
Algorithm 2 PseudoDescend
procedure PseudoDescend($E(\Lambda),c,m,u,r,\gamma$)
     Visit $\Lambda$
     $\tilde{r}\leftarrow r$
     $\tilde{E}\leftarrow E$
     for $\sigma$ from $c+1$ to $c+u-1$ $\triangleright$ where right generators are strong
         if $\sigma\neq 2m$ then
              Descend($\tilde{E},c-1,\sigma+1,m,u,\tilde{r},\gamma$)
              $\tilde{r}\leftarrow\tilde{r}-1$
         end if
     end for
     for $\sigma$ from $c+u$ to $c+m-1$ $\triangleright$ where right generators are weak
         if $\sigma\neq 2m$ then
              $\tilde{r}\leftarrow\tilde{r}-1$
              Descend($\tilde{E},c-1,\sigma+1,m,u,\tilde{r},\gamma$)
         end if
     end for
end procedure
Now we are ready to define a general exploring algorithm. For an integer $m$ let $O_{m}$ be the unique ordinary semigroup of multiplicity $m$. For $2\leqslant u\leqslant m$, let $P_{m,u}$ be the unique pseudo-ordinary semigroup with multiplicity $m$ and jump $u$. For $m+1\leqslant F\leqslant 2m-1$, let $Q_{m,F}$ be the unique quasi-ordinary semigroup with multiplicity $m$ and Frobenius number $F$. Let $H_{g}$ be the hyperelliptic semigroup of genus $g$, that is, the semigroup generated by $2$ and $12g+$ 1.
Let ${\mathscr{T}}_{m}$ be the subtree of ${\mathscr{T}}$ with all the semigroups of multiplicity $m$. It is the subtree that contains $O_{m}$ together with all its descendants, except for the branch emerging from its unique ordinary child, $O_{m+1}$.
In turn, ${\mathscr{T}}_{m}$ can be splitted into its $m-1$ subtrees ${\mathscr{T}}_{m,u}$, for $2\leqslant u\leqslant m$, each of which contains all the semigroups of multiplicity $m$ and jump $u$ and the tree ${\mathscr{Q}}_{m}$ which is rooted in $O_{m}$ and contains all the quasi-ordinary semigroups of multiplicity $m$ and all their respective descendants. That is, ${\mathscr{Q}}_{m}$ contains all the semigroups with multiplicity $m$ and jump $u=1$.
Notice that ${\mathscr{T}}_{m,u}$ is the subtree of ${\mathscr{T}}_{m}$ that contains $P_{m,u}$ together with all its descendants, except for the branch emerging from its unique pseudo-ordinary child, $P_{m,u+1}$, should $u<m$.
Combining the previous procedures Descend and PseudoDescend we can define the algorithm ExploreTree detailed in Algorithm 3 for exploring ${\mathscr{T}}$ up to a given genus $\gamma$. It separately explores the trees ${\mathscr{T}}_{1},\dots,{\mathscr{T}}_{\gamma+1}$, and, within the exploration of each ${\mathscr{T}}_{m}$, it separately explores the tree ${\mathscr{Q}}_{m}$ and the trees ${\mathscr{T}}_{m,2}$, ${\mathscr{T}}_{m,3}$, …, ${\mathscr{T}}_{m,min}$, where $min=\min\{m,\gamma+2-m\}$.
Hence, ExploreTree can be parallelized by the multiplicity and the jump in a straighforward way. In turn, the exploration of ${\mathscr{Q}}_{m}$ can be parallelized by the $m$’th gap, that is, the Frobenius number.
Algorithm 3 ExploreTree
procedure ExploreTree($\gamma$)
     Visit ${\mathbb{N}}_{0}$triangleright The unique semigroup with $m=1$
     for $g$ from $1$ to $\gamma$ do $\triangleright$ The semigroups with $m=2$
         Visit $H_{g}$
     end for
     for $m$ from $3$ to $\gamma$ do
         Visit $O_{m}$
         $min\leftarrow\min\{m,\gamma+2-m\}$
         for $u$ from 2 to $min-1$ do
              Visit $P_{m,u}$
              PseudoDescend($E(P_{m,u}),m+u,m,u,m-2,\gamma$) $\triangleright$ $c=m+u$, $r=m-2$
         end for
         Visit $P_{m,min}$
         if $min<\gamma+2-m$ then
              PseudoDescend($E(P_{m,m}),2m,m,m,m-1,\gamma$) $\triangleright$ $u=m$, $r=m-1$
         end if
         $r\leftarrow m-3$
         for $\sigma$ from $m+2$ to $22m-$ 2 do $\triangleright$ loop on the quasi-ordinaries
              Descend($E(Q_{m,\sigma}),m,\sigma+1,m,1,r,\gamma$)
              $r\leftarrow r-1$
         end for
         Visit $Q_{m,2m-1}$triangleright $Q_{m,2m-1}$ has no descendants if $m>2$
     end for
     Visit $O_{\gamma+1}$triangleright The unique semigroup with $g\leqslant\gamma$ and $m=\gamma+1$
end procedure
3 Encoding by shrinking
3.1 A new encoding
Given a numerical semigroup $\Lambda$ denote $L(\Lambda)$ its set of left elements, that is, its elements that are smaller than its Frobenius number. Now, given a numerical semigroup $\Lambda$, define
$\omega(\Lambda)$ ::=:= $\gcd(L(\Lambda))$
$\overline{\Lambda}$ ::=:= $\left\langle\frac{L(\Lambda)}{\omega}\right\rangle$
The tuple $\omega(\Lambda),\overline{\Lambda}$ is an encoding of $\Lambda$ as defined in Section 2. Indeed, together with $c(\Lambda)$ it uniquely determines $\Lambda$, since
$\Lambda=\omega(\Lambda)\overline{\Lambda}\cup\{c(\Lambda)+{\mathbb{N}}_{0}\}.$
In the definition of an encoding system we required its parameters to be finite. Here we assume that $\overline{\Lambda}$ is represented by a finite set, for instance, by its left elements. We call $\overline{\Lambda}$ the shrinking of $\Lambda$.
3.2 CheckRightGenerator and CheckStrongGenerator
Next we show that we have a direct way to determine from $\overline{\Lambda}$ and $\omega(\Lambda)$ the right generators of $\Lambda$, and that it is equaly easy to identify the strong generators.
Lemma 3.1.
Let $\Lambda$ be a numerical semigroup and let $c=c(\Lambda)$, $m=m(\Lambda)$, $u=u(\Lambda)$, $\omega=\omega(\Lambda)$.
  1. 1.
    Let $c\leqslant\sigma<c+u$. The element $\sigma\in\Lambda$ is a right generator of $\Lambda$ if and only if either
    1. (a)
      $\sigma\not\equiv 0\mod\omega$
    2. (b)
      $\sigma\equiv 0\mod\omega$ and $\frac{\sigma}{\omega}\not\in\overline{\Lambda}$.
  2. 2.
    In case (a), $\sigma$ is strong if and only if $\sigma<c+u$.
  3. 3.
    In case (b), $\sigma$ is strong if and only if $\frac{\sigma+m}{\omega(\Lambda\setminus\{\sigma\})}\not\in\overline{\Lambda\setminus\{\sigma\}}$.
Proof.
It is obvious that if $\sigma\not\equiv 0\mod\omega$ then $\sigma$ is a right generator. If $\sigma\equiv 0\mod\omega$, then it is not a right generator if and only if it is generated by the left elements, which is equivalent to $\frac{\sigma}{\omega}\in\overline{\Lambda}$.
For the second statement, on one hand it follows from Lemma 2.1 that if $\sigma\geqslant c+u$ then $\sigma$ is not strong. Now, suppose that $\sigma<c+u$ and suppose that $\sigma$ is not strong, that is, $\sigma+m=a+b$ with $\{a,b\}\neq\{\sigma,m\}$ and $\{a,b\}\neq\{0,\sigma+m\}$. Since $a,b\neq 0,m$, then $a,b\geqslant m+u>m+\sigma-c$ and, by the equality $a+b=\sigma+m$ we deduce that $\left\{\begin{array}[]{l}\sigma+m>a+m+\sigma-c\\ \sigma+m>b+m+\sigma-c\end{array}\right.$ and, so, $a,b<c$. Then, $a,b\equiv 0\mod\omega$ and $a+b\equiv 0\mod\omega$ implying that $\sigma\equiv 0\mod\omega$, a contradiction.
For the third item, we know that $\sigma$ is strong if and only if $\sigma+m$ is a minimal generator of $\Lambda\setminus\{\sigma\}$, and, by the first item, this is equivalent to $\frac{\sigma+m}{\omega(\Lambda\setminus\{\sigma\})}\not\in\overline{\Lambda\setminus\{\sigma\}}$. ∎
3.3 Encoding pseudo-ordinary and quasi-ordinary semigroups and EncodingFromParent and EncodingFromPredecessorSibling
Semigroups generated by an interval
In order to explain the encoding procedures we need some results on semiroups generated by intervals. Let $\Lambda_{\{i,\dots,j\}}$ be the semigroup generated by the interval $\{i,i+1,\dots,j\}$. Notice that it is the union of sets $S_{k}=\{ki,\dots,kj\}$.
Lemma 3.2.
The conductor of $\Lambda_{\{i,\dots,j\}}$ is $i\lfloor\frac{j-2}{j-i}\rfloor$. The genus of $\Lambda_{\{i,\dots,j\}}$ is $\sum_{k=1}^{i\lfloor\frac{j-2}{j-i}\rfloor}(i+(k-1)(i-j)-1)$.
Proof.
Let $D_{1}$ be the set of gaps between 00 and $S_{1}$ and, in general, $D_{k}$ the set of gaps between $S_{k-1}$ and $S_{k}$. One can check that #$\#D_{k}=i+(k-1)(i-j)-1$ as far as this is a positive amount, although we will also extend this equality for the case of negative cardinality. Let $k_{c}$ be such that the conductor of $\Lambda$ is the smallest element in $S_{k_{c}}$. Equivalently,
$k_{c}=\frac{c}{i}.$ (1)
Notice that the Frobenius number of $\Lambda_{\{i,\dots,j\}}$ is the largest element of $D_{k_{c}}$ and $D_{i}\neq 0$ if and only if $i\leqslant k_{c}$. Hence, $k_{c}$ is the largest element such that #$\#D_{k}\geqslant 1$, that is,
$i+(k_{c}-1)(i-j)-1$$\geqslant$ 11 11
$i+(k_{c}-1)(i-j)$$\geqslant$ 22 22
$k_{c}$$\leqslant$ $\frac{2-i}{i-j}+1=\frac{2-j}{i-j}=\frac{j-2}{j-i}$
So,
$k_{c}=\left\lfloor\frac{j-2}{j-i}\right\rfloor.$
Now the result follows from (1). ∎
Encoding pseudo-ordinary and quasi-ordinary semigroups
It is easy to check that $\omega(P_{m,u})=m$ and $\overline{P}_{m,u}={\mathbb{N}}_{0}$. In particular $Q_{m,m+1}=P_{m,2}$ and so $\omega(Q_{m,m+1})=m$ and $\overline{Q}_{m,m+1}={\mathbb{N}}_{0}$. Similarly, if $F>m+1$, then $L(Q_{m,F})=\{m,m+1,\dots,F-1\}$ and, so, $\omega(Q_{m,F})=1$ and $\overline{Q}_{m,F}=\Lambda_{\{m,\dots,F-1\}}$.
Encoding from the parent
Let us see how we can encode a semigroup from the encoding of its parent. If $A,B$ are two sets, we will use $A+B$ to refer to the set $\{a+b:a\in A,b\in B\}$. If $\Lambda_{1},\Lambda_{2}$ are submonoids of ${\mathbb{N}}_{0}$, then $\Lambda_{1}+\Lambda_{2}$ is also a submonoid and, if either $\Lambda_{1}$ or $\Lambda_{2}$ is a numerical semigroup, then $\Lambda_{1}+\Lambda_{2}$ is a numerical semigroup and, indeed, it is the minimum semigroup that contains $\Lambda_{1}$ and $\Lambda_{2}$.
Lemma 3.3.
Let $\Lambda$ be a numerical semigroup with conductor $c$. Let $\omega=\omega(\Lambda)$, $\tilde{\Lambda}=\Lambda\setminus\{\sigma\}$ and $\tilde{\omega}=\omega(\tilde{\Lambda})$.
  • If $\sigma=c$,
    • $\tilde{\omega}=\omega$,
    • $\overline{\tilde{\Lambda}}=\overline{\Lambda}$,
    • $c(\overline{\tilde{\Lambda}})=c(\overline{\Lambda})$.
  • If $\sigma=c+1$,
    • $\tilde{\omega}=\gcd(\omega,c)$,
    • $\overline{\tilde{\Lambda}}=\frac{\omega}{\tilde{\omega}}\overline{\Lambda}+c{\mathbb{N}}_{0}$,
    • $c(\overline{\tilde{\Lambda}})\leqslant c(\overline{\Lambda})\frac{\omega}{\tilde{\omega}}+(\frac{c}{\tilde{\omega}}-1)(\frac{\omega}{\tilde{\omega}}-1)$.
  • If $\sigma>c+1$,
    • $\tilde{\omega}=1$,
    • $\overline{\tilde{\Lambda}}=\omega\overline{\Lambda}+\Lambda_{c,\dots,{\sigma-1}}$,
    • $c(\overline{\tilde{\Lambda}})\leqslant c(\overline{\Lambda})\omega+(\lfloor\frac{\omega-2}{\sigma-c-1}\rfloor+1)c$
Proof.
The case $\sigma=c$ is clear.
Suppose $\sigma=c+1$. Now $\overline{\tilde{\Lambda}}$ contains $c(\overline{\Lambda})\frac{\omega}{\tilde{\omega}}$ and $\frac{c}{\tilde{\omega}}$, which are coprime. Now, the list $c(\overline{\Lambda})\frac{\omega}{\tilde{\omega}}$, $c(\overline{\Lambda})\frac{\omega}{\tilde{\omega}}+\frac{c}{\tilde{\omega}}$, $c(\overline{\Lambda})\frac{\omega}{\tilde{\omega}}+2\frac{c}{\tilde{\omega}}$, $c(\overline{\Lambda})\frac{\omega}{\tilde{\omega}}+3\frac{c}{\tilde{\omega}}$, …, $c(\overline{\Lambda})\frac{\omega}{\tilde{\omega}}+(\frac{\omega}{\tilde{\omega}}-1)\frac{c}{\tilde{\omega}}$ contains $\frac{\omega}{\tilde{\omega}}$ consecutive elements of $\overline{\tilde{\Lambda}}$ which are not congruent modulo $\frac{\omega}{\tilde{\omega}}$. But, since $(c(\overline{\Lambda})+i)\frac{\omega}{\tilde{\omega}}$ also belongs to $\overline{\tilde{\Lambda}}$ for any positive integer $i$, we can add any multiple of $\frac{\omega}{\tilde{\omega}}$ to any element of the previous list and still obtain elements in $\overline{\tilde{\Lambda}}$. Finally, it can be seen that the elements between $c(\overline{\Lambda})\frac{\omega}{\tilde{\omega}}+(\frac{\omega}{\tilde{\omega}}-1)\frac{c}{\tilde{\omega}}-\frac{\omega}{\tilde{\omega}}+1$ and $c(\overline{\Lambda})\frac{\omega}{\tilde{\omega}}+(\frac{\omega}{\tilde{\omega}}-1)\frac{c}{\tilde{\omega}}$ all belong to $\overline{\tilde{\Lambda}}$ and, so, $c(\overline{\tilde{\Lambda}})\leqslant c(\overline{\Lambda})\frac{\omega}{\tilde{\omega}}+(\frac{\omega}{\tilde{\omega}}-1)\frac{c}{\tilde{\omega}}-\frac{\omega}{\tilde{\omega}}+1=c(\overline{\Lambda})+(\frac{\omega}{\tilde{\omega}}-1)(\frac{c}{\tilde{\omega}}-1)$.
Suppose $\sigma>c+1$. It is obvious that $\tilde{\omega}=1$. The semigroup $\overline{\tilde{\Lambda}}$ must contain $\Lambda_{c,\dots,\sigma-1}$ and, so, it contains the union of sets $S_{k}=\{kc,\dots,k(\sigma-1)\}$, which have $k(\sigma-1-c)+1$ consecutive elements. In particular, if $\ell=\lceil\frac{\omega-1}{\sigma-1-c}\rceil$, the set $S_{\ell}$ contains $\omega$ consecutive elements. From this we deduce that the intervals $(c(\overline{\Lambda})+i)\omega+S_{\ell}$ all contain $\omega$ consecutive elements and they cover all the integers greater than or equal to $c(\overline{\Lambda})+\ell c$. Hence, the conductor of $\overline{\tilde{\Lambda}}$ is at most $c(\overline{\Lambda})\omega+\ell c$ and the result follows. ∎
Encoding from the predecessor sibling
It is now straightforward proving the following lemma.
Lemma 3.4.
Let $\Lambda$ be a numerical semigroup with conductor $c$. Suppose that the right generators of $\Lambda$ are $\sigma_{1}<\dots<\sigma_{r}$. Suppose that $2\leqslant i\leqslant r$ and that $\sigma_{i-1}\neq c$. Let $\tilde{\Lambda}=\Lambda\setminus\{\sigma_{i}\}$, and let $\tilde{\Lambda}^{\prime}=\Lambda\setminus\{\sigma_{i-1}\}$. If $\omega(\tilde{\Lambda}^{\prime})=1$, then
    • $\tilde{\omega}=1$,
    • $\overline{\tilde{\Lambda}}={{\overline{\tilde{\Lambda}}}^{\prime}}+\Lambda_{\sigma_{i-1},\dots,{\sigma_{i}-1}}$,
    • $c(\overline{\tilde{\Lambda}})\leqslant c({\overline{\tilde{\Lambda}}}^{\prime})$
Notice that the condition $\omega(\tilde{\Lambda}^{\prime})=1$ is satisfied whenever $\sigma_{i-1}\geqslant n+3$.
4 The unleaved tree
The interest of encoding by the gcd and shrinking is the next result, which was first proved in [7], although in another context and with different notation.
Lemma 4.1.
[7, Theorem 10] If $\omega(\Lambda)\neq 1$, then $\Lambda$ has descendants of any given genus. If $\omega(\Lambda)=1$, the maximum genus of the descendants of $\Lambda$ is the genus of $\overline{\Lambda}$ and there is only one descendant of that genus.
Hence, if we just want to visit the semigroups of genus $\gamma$, we can trim any numerical semigroup $\Lambda$ of $\mathscr{T}$ satisfying both that $\omega(\Lambda)=1$ and that the genus of $\overline{\Lambda}$ is smaller than $\gamma$, together with all its descedants. We call unleaved tree of genus $\gamma$ the subtree that we obtain. This tree has no leaves of genus smaller than $\gamma$ and all its leaves are exactly the semigroups of genus $\gamma$. In contrast, we will call complete tree of genus $\gamma$ the subtree of ${\mathscr{T}}$ that contains all numerical semigroups of genus up to $\gamma$.
In Figure 1 we represented the complete tree of genus $9$ (left), the unleaved tree of genus $9$ (center), and the edges of the complete tree not appearing in the unleaved tree (right).
Figure 1: Complete tree, unleaved tree, and their difference for genus $\gamma=9$
If our objective in the exploration of the tree is counting, we can also avoid visiting the nodes such that the genus of $\overline{\Lambda}$ is exactly equal to $\gamma$ and just count $1$.
A fact that makes trimming very efficacious is that it is proved that most numerical semigroups belong to finite chains [10], that is, most numerical semigroups have the gcd of their left elements equal to $1$.
Before deciding whether we trim a semigroup and all its descendants we need to encode it, which takes some computing time. Thus, an important trick of our algorithm is not encoding nodes that we know a priori that will be trimmed. The result in next lemma is in this direction.
Lemma 4.2.
If $\Lambda$, $\Lambda^{\prime}$ are siblings in the semigroup tree with $F(\Lambda^{\prime})<F(\Lambda)$, then if $\Lambda^{\prime}$ has no descendants of genus $\gamma$ then neither does $\Lambda$.
Proof.
If $\Lambda^{\prime}$ has no descendants then $\omega(\Lambda^{\prime})=1$. The left elements of $\Lambda$ are the left elements of $\Lambda^{\prime}$ together with the interval $\{F(\Lambda^{\prime}),\dots,F(\Lambda)-1\}$, which is not empty and which contains elements not in $\Lambda^{\prime}$. So, $\omega(\Lambda)=\omega(\Lambda^{\prime})=1$ and $\overline{\Lambda^{\prime}}$ is strictly contained in $\overline{\Lambda}$. So, $g(\overline{\Lambda^{\prime}})>g(\overline{\Lambda})$. Hence, if $g(\overline{\Lambda^{\prime}})<\gamma$, then $g(\overline{\Lambda})<\gamma$. ∎
All these results suggest the function DescendAndTrim shown in Algorithm 4 in substitution of Descend, when the exploration of the tree is aimed at counting semigroups of genus $\gamma$. If the exploration of the tree is aimed at visiting all nodes of genus $\gamma$, and not just counting, then a statement for visiting $\Lambda$ should be placed at the beginning, all the lines with the parameter count, as well as all lines from line 59 to the end should be omitted, and line 5 should be replaced by “if $g<\gamma$ then”.
Notice that the procedure PseudoDescend can be equally optimized using the same idea of trimming. We call PseudoDescendAndTrim the new function. For the sake of brevity we do not put here the associated pseudo-code.
Algorithm 4 DescendAndTrim
1:function DescendAndTrim($\omega,\overline{\Lambda},g,c,m,u,r,\gamma$)
2:   count $\leftarrow 0$
3:   $\tilde{r}\leftarrow r$
4:   $\tilde{\omega},\overline{\tilde{\Lambda}}\leftarrow\omega,\overline{\Lambda}$
5:   if $g<\gamma-2$ then
6:     keepgoing $\leftarrow$ true
7:     for $\sigma$ from $c$ to $c+u-1$, while keepgoing  do
8:       if $\sigma\not\equiv 0\mod\omega$ or $\frac{\sigma}{\omega}\not\in\overline{\Lambda}$  then
9:         if $\sigma$ is primogenial or the predecessor sibling is $c$ then
10:            $\tilde{\omega},\overline{\tilde{\Lambda}}\leftarrow$ EFP($\omega,\overline{\Lambda},g,c,m,u,\tilde{r},\sigma$) $\triangleright$ Lemma 3.3
11:         else
12:            $\tilde{\omega},\overline{\tilde{\Lambda}}\leftarrow$ EFPS($\tilde{\omega},\overline{\tilde{\Lambda}},g,c,m,u,\tilde{r},\sigma$) $\triangleright$ Lemma 3.4
13:         end if
14:         if $\tilde{\omega}\neq 1$ then
15:            if $\sigma\not\equiv 0\mod\tilde{\omega}$ or $\frac{\sigma+m}{\tilde{\omega}}\not\in\overline{\tilde{\Lambda}}$ then
16:              count $\leftarrow$ count$+$DescendAndTrim($\tilde{\omega},\overline{\tilde{\Lambda}},g+1,\sigma+1,m,u,\tilde{r},\gamma$)
17:              $\tilde{r}\leftarrow\tilde{r}-1$
18:            else
19:              $\tilde{r}\leftarrow\tilde{r}-1$
20:              count $\leftarrow$ count$+$DescendAndTrim($\tilde{\omega},\overline{\tilde{\Lambda}},g+1,\sigma+1,m,u,\tilde{r},\gamma$)
21:            end if
22:         else
23:            if genus$(\overline{\tilde{\Lambda}})\leqslant\gamma$ then
24:              if genus$(\overline{\tilde{\Lambda}})=\gamma$ then
25:                count $\leftarrow$ count$+1$
26:              end if
27:              keepgoing $\leftarrow$ false
28:            else
29:              if $\sigma+m\not\in\overline{\tilde{\Lambda}}$ then
30:                count $\leftarrow$ count$+$DescendAndTrim($\tilde{\omega},\overline{\tilde{\Lambda}},g+1,\sigma+1,m,u,\tilde{r},\gamma$)
31:                $\tilde{r}\leftarrow\tilde{r}-1$
32:              else
33:                $\tilde{r}\leftarrow\tilde{r}-1$
34:                count $\leftarrow$ count$+$DescendAndTrim($\tilde{\omega},\overline{\tilde{\Lambda}},g+1,\sigma+1,m,u,\tilde{r},\gamma$)
35:              end if
36:            end if
37:         end if
38:       end if
39:     end for
40:     while keepgoing and $\tilde{r}>1$ do
41$\triangleright$ no strong generators here, so if $\tilde{r}=1$ there will be no grand-children
42:       if $\sigma\not\equiv 0\mod\omega$ or $\frac{\sigma}{\omega}\not\in\overline{\Lambda}$  then
43:         if $\sigma$ is primogenial or the predecessor sibling is $c$ then
44:            $\tilde{\omega},\tilde{\Lambda}\leftarrow$ EFP($\tilde{\omega},\overline{\tilde{\Lambda}},g,c,m,u,r,\sigma$) $\triangleright$ Lemma 3.3
45:         else
46:            $\tilde{\omega},\tilde{\Lambda}\leftarrow$ EFPS($\tilde{\omega},\overline{\tilde{\Lambda}},g,c,m,u,r,\sigma$) $\triangleright$ Lemma 3.4
47:         end if
48:         if genus$(\overline{\tilde{\Lambda}})\leqslant\gamma$ then
49:            if genus$(\overline{\tilde{\Lambda}})=\gamma$ then
50:              count $\leftarrow$ count$+1$
51:            end if
52:            keepgoing $\leftarrow$ false
53:         else
54:            $\tilde{r}\leftarrow\tilde{r}-1$
55:            count $\leftarrow$ count$+$DescendAndTrim($\tilde{\omega},\overline{\tilde{\Lambda}},g+1,\sigma+1,m,u,\tilde{r},\gamma$)
56:         end if
57:       end if
58:     end while
59:   $\triangleright$ if $g=\gamma-2$
60:     for $\sigma$ from $c$ to $c+u-1$ do
61:       if $\sigma\not\equiv 0\mod\omega$ or $\frac{\sigma}{\omega}\not\in\overline{\Lambda}$  then
62:         if $\sigma$ is primogenial or the predecessor sibling is $c$ then
63:            $\tilde{\omega},\tilde{\Lambda}\leftarrow$ EFP($\tilde{\omega},\overline{\tilde{\Lambda}},g,c,m,u,r,\sigma$) $\triangleright$ Lemma 3.3
64:         else
65:            $\tilde{\omega},\tilde{\Lambda}\leftarrow$ EFPS($\tilde{\omega},\overline{\tilde{\Lambda}},g,c,m,u,r,\sigma$) $\triangleright$ Lemma 3.4
66:         end if
67:         if $\sigma\not\equiv 0\mod\tilde{\omega}$ or $\frac{\sigma+m}{\tilde{\omega}}\not\in\overline{\tilde{\Lambda}}$ then
68:            count $\leftarrow$ count$+r$
69:            $\tilde{r}\leftarrow\tilde{r}-1$
70:         else
71:            $\tilde{r}\leftarrow\tilde{r}-1$
72:            count $\leftarrow$ count$+r$
73:         end if
74:       end if
75:     end for
76:     while $\tilde{r}>1$ do
77:       $\tilde{r}\leftarrow\tilde{r}-1$
78:       count $\leftarrow$ count$+r$
79:     end while
80:   end if
81:   return count
82:end function
5 More unvisited nodes
Lemma 5.1.
[9, Section 5] The number of children of $P_{m,u}$ is $m-1$.
The number of grandchildren of $P_{m,u}$ is
$\left\{\begin{array}[]{ll}\binom{m-1}{2}+u&\mbox{ if }2u\leqslant m\\ \binom{m-1}{2}+u-1&\mbox{ otherwise}\\ \end{array}\right.$
Notice that $P_{m,u^{\prime}}$ is a descendant of $P_{m,u}$ as far as $u^{\prime}>u$. If we only want to count the semigroups of genus $\gamma$, in our algorithm we need only to descend the pseudo-ordinary semigroups $P_{m,u}$ for $u\leqslant\gamma-m-1$, as far as $\gamma-m\leqslant m$. Indeed, $P_{m,\gamma-m+1}$ and $P_{m,\gamma-m+1}$ (should they exist) are descendants of $P_{m,\gamma-m}$, and the descendants of $P_{m,\gamma-m}$ of genus $\gamma$ are exactly its grandchildren, which can be counted by the formula in Lemma 5.1, without needing to visit them.
From another perspective, Rosales proved the result in the next lemma in [15].
Lemma 5.2.
[15, Corollary 10] There is exactly one semigroup of genus $\gamma$ and multiplicity $2$ and there are $\gamma-\lfloor\frac{2\gamma-1}{3}\rfloor$ semigroups of genus $g$ and multiplicity $3$
Lemma 5.3.
If $\gamma\geqslant 8$, then there are $\binom{\gamma-4}{4}+\binom{\gamma-2}{3}+\binom{\gamma-5}{2}+6\gamma-14$ numerical semigroups of genus $\gamma$ and multiplicity larger than or equal to $\gamma-3$.
Proof.
It was proved in [6, Theorem 9] that for $m\geqslant 4$, the ordinary semigroup of multiplicity $m$ has exactly $\binom{m}{3}+3m+3$ great-grandchildren. This implies that the ordinary semigroup of multiplicity $m=\gamma-2$ has $\binom{\gamma-2}{3}+3\gamma-3$ descendants of genus $\gamma$, and, so, there are $\binom{\gamma-2}{3}+3\gamma-3$ semigroups of genus $\gamma$ and multiplicity larger than or equal to $\gamma-2$.
Now we claim that there are excatly $\binom{\gamma-4}{4}+\binom{\gamma-5}{2}+3\gamma-11$ semigroups of genus $g$ and multiplicity exactly equal to $\gamma-3$.
Suppose that a semigroup has genus $g$ and multiplicity $m=\gamma-3$. In particular, the semigroup has exactly four gaps larger than the multiplicity. Define $A_{1}$ as the set of congruence classes modulo $m$ of its gaps between $m$ and $2m$ and define subsequently $A_{i}$ as the set of classes modulo $m$ of its gaps between $i\,m$ and $(i+1)m$. It is obvious that $A_{i+1}\subseteq A_{i}$. Let $j$ be the maximum integer such that $A_{j}\neq\emptyset$. The genus of the semigroup is $m+3$ and the Frobenius number is $j\,m+\max(A_{j})$, from which $j\,m+\max(A_{j})\leqslant 2m+5$ and, so $51\leqslant\max(A_{j})\leqslant(2-j)m+$ 5. This implies that $j\leqslant 2+\frac{4}{m}$ and so, for $m>4$ (which is the case if $m=\gamma-3$ and $\gamma\geqslant 8$), we have $j\leqslant 2$. In particular, #$\#A_{1}+\#A_{2}=4$ and, if $A_{2}\neq\emptyset$, then $\max(A_{2})\leqslant 5$.
Consider all the possible cases dependning on the cardinality of $A_{1}$. If #$\#A_{1}=4$, then there are no restrictions on the gaps among the $m-1$ possibilities between $m+1$ and $12m-$ 1. Hence, there are $\binom{m-1}{4}=\binom{\gamma-4}{4}$ options. If #$\#A_{1}=3$, then #$\#A_{2}=1$ and, since $\max(A_{2})\leqslant 5$, there are only five options: If $A_{2}=\{1\}$, then $A_{1}$ can be any subset of three elements between $1$ and $m-1$ containing $1$. There are $\binom{m-2}{2}=\binom{\gamma-5}{2}$ such options; If $A_{2}=\{2\}$, then $A_{1}$ can be any subset of three elements containing $\{1,2\}$. There are $m-3=\gamma-6$ options; If $A_{2}=\{3\}$, then $A_{1}$ can be any subset of three elements either containing $\{1,3\}$ or containing $\{2,3\}$. There are $132(m-3)-1=2\gamma-$ 13 options; If $A_{2}=\{4\}$, then $A_{1}$ is either $\{1,2,4\}$ or $\{2,3,4\}$. Those are $2$ options; If $A_{2}=\{5\}$, then $A_{1}$ is either $\{1,2,5\},\{1,3,5\},\{2,4,5\},\{3,4,5\}$. Those are $4$ options. If #$\#A_{1}=2$, then #$\#A_{2}=2$ and there are only two options: either $A_{1}=A_{2}=\{1,2\}$ or $A_{1}=A_{2}=\{1,3\}$.
We conclude that $n_{\gamma,\gamma-3}=\binom{\gamma-4}{4}+\binom{\gamma-5}{2}+\gamma-3\gamma-11$ and the result of the lemma follows. ∎
6 The algorithm ExploreUnleavedTree
All these results give rise to the trimming version of the algorithm ExploreTree when used to count semigroups of genus $\gamma$, for $\gamma\geqslant 8$, without needing to visit them. It is the algorithm ExploreUnleavedTree shown in Algorithm 5.
Algorithm 5 ExploreUnleavedTree
function ExploreUnleavedTree($\gamma$)
    count $\leftarrow\binom{\gamma-4}{4}+\binom{\gamma-2}{3}+\binom{\gamma-5}{2}-\lfloor\frac{2\gamma-1}{3}\rfloor+7\gamma-13$$\triangleright$ Lemma 5.2 and Lemma 5.3
    for $m$ from $4$ to $\gamma-4$ do
         $min\leftarrow\min\{m,\gamma-m\}$
         for $u$ from 2 to $min-1$ do
             count $\leftarrow$ count$+$PseudoDescendAndTrim($m,{\mathbb{N}}_{0},m+u,m,u,m-2,\gamma$) $\triangleright$ $\omega(P_{m,u})=m$, $\overline{P_{m,u}}={\mathbb{N}_{0}}$
         end for
         if $min<\gamma-m$ then
             count $\leftarrow$ count$+$PseudoDescendAndTrim($m,{\mathbb{N}}_{0},2m,m,m,m-1,\gamma$) $\triangleright$ $\omega(P_{m,m})=m$, $\overline{P_{m,m}}={\mathbb{N}_{0}}$
         else
             count $\leftarrow$ count$+\gamma-m+\binom{m-1}{2}$
             if $2\gamma>3m$ then
                 count $\leftarrow$ count11-1- 1
             end if
         end if
         $r\leftarrow m-3$
         for $\sigma$ from $m+2$ to $22m-$ 2 while genus$(\Lambda_{m,\dots,\sigma-1})\geqslant\gamma$ do
             if genus$(\Lambda_{m,\dots,\sigma-1})=\gamma$ then
                 count $\leftarrow$ count$+1$
             else
[…]
                 $r\leftarrow r-1$
             end if
         end for
    end for
    return count
end function
We now want to have a look at the number of numerical semigroups that are encoded in the algorithm ExploreUnleavedTree. Note that there are semigroups which are encoded and then trimmed after checking that they have no descendants of genus $\gamma$. On the other hand, there are semigroups which belong to the unleaved tree but need not to be encoded, as for instance the semigroups of smallest or largest possible multiplicities, the pseudo-ordinary semigroups, or the semigroups whose descendants in ${\mathscr{T}}$ have maximum dept equal to $\gamma$, because we already know how many descendants they have of genus $\gamma$.
In Figure 2 we represented the complete tree of genus $9$ (left), the subset of those semigroups that are encoded by the algorithm (center), and the subset of those semigroups that are not encoded (right).
Figure 2: All semigroups of genus up to $9$; subset of semigroups encoded by the algorithm; subset of non encoded semigroups
The number of encoded nodes is expected to be proportional to the computation time. In Table 1 we give the number of nodes in the complete tree, number of nodes in the unleaved tree and the number of nodes encoded by the algorithm ExploreUnleavedTree for $\gamma$ from $10$ to $40$. We can see that the number of nodes in the unleaved tree is around $63\%$ of the number of nodes of the complete tree. Also, the number of semigroups encoded by our algorithm is just around $15\%$ of the number of nodes of the complete tree, and it is also less than half $n_{g}$. This is the main reason for our algorithm to go much faster than any algorithm exploring all nodes of ${\mathscr{T}}_{g}$.
Table 1: For each genus $\gamma$, number $n_{\gamma}$ of numerical semigroups of genus $\gamma$, number of nodes in the complete tree of genus $\gamma$, number of nodes in the unleaved tree of genus $\gamma$, number of nodes encoded by the algorithm for genus $\gamma$, and their respective percentages with respect to the number of nodes in the complete tree.
$g$ 10 15 20 25 30 35 40
$n_{g}$ 204 2857 37396 467224 5646773 66687201 774614284
complete tree 478 6964 93142 1179597 14396338 171202690 1998799015
unleaved tree 364 4833 61469 759972 9146174 107815637 1251716100
(wrt. complete) ($76\%$) ($69\%$) ($66\%$) ($64\%$) ($64\%$) ($63\%$) ($63\%$)
encoded nodes 61 1325 16774 196433 2282567 26454236 304794995
(wrt. complete) ($13\%$) ($19\%$) ($18\%$) ($17\%$) ($16\%$) ($15\%$) ($15\%$)
Acknowledgement
This research is supported by the project “HERMES” funded by the European Union NextGenerationEU/PRTR via INCIBE, by project PID2021-124928NB-I00 “ACITHEC” funded by MCIN/AEI/10.13039/501100011033/ FEDER, UE, and by the grant 2021SGR 00115 from the Government of Catalonia.
The author is grateful to Christopher O’Neill for very interesting discussions and for suggesting the formula for $n_{\gamma,\gamma-3}$ used in the proof of Lemma 5.3. She is also grateful to Armando Martín López and Jose Luis Muñoz for facilitating the computational resources that made it possible to compute $n_{76}$ and $n_{77}$.